OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 10x2 + 24x ? –6, 0, 4 6, 0, –4 –6, 0, –4 6, 0, 4

OpenStudy (hoblos):

x3 + 10x2 + 24x = x(x2+10x+24) = x(x+4)(x+6) so the zeros are –6, 0, –4

OpenStudy (anonymous):

I forgot about the zero

OpenStudy (anonymous):

You are right -6, 0, -4